OpenStudy (anonymous):

Use basic identities to simplify the expression: cos(x)-cos^3(x)

OpenStudy (anonymous):

I have no clue which equations to even look at. I know the answer but not how to get to it.

OpenStudy (anonymous):

factorise cos x

OpenStudy (anonymous):

I dont know how :/

OpenStudy (anonymous):

can you factor this \[a-a^3\]

OpenStudy (anonymous):

factor a

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

I dont understand...its in its lowest form?

OpenStudy (anonymous):

\[a-a^3=a(1-a^2)\]

OpenStudy (anonymous):

do you understand this

OpenStudy (anonymous):

wouldnt it be a(1-a)? cause youre gonna factor?

OpenStudy (anonymous):

but it is a^2 cos when we multiply out that it has to give us the original expression ie a-a^3

OpenStudy (anonymous):

but a(1-a)=a-a*a=a-a^2 not a-a^3

OpenStudy (anonymous):

I dont follow :(

OpenStudy (anonymous):

can you factor this \[x^3+x^4\]

OpenStudy (anonymous):

is that a superscript 3 or 2

OpenStudy (anonymous):

x^3(1+x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

so \[x-x^3\]

OpenStudy (anonymous):

factor this

OpenStudy (anonymous):

cosx(1-cos^2x)

OpenStudy (anonymous):

yes so use this identity \[\cos ^2x+\sin^2x=1\]...hint take \[\cos ^2x\] to the other side

OpenStudy (anonymous):

sin^2x-1=cos^2x?

OpenStudy (anonymous):

or do i leave the 1

OpenStudy (anonymous):

yes leave the 1

OpenStudy (anonymous):

sin^2x=cos^2x+1

OpenStudy (anonymous):

cos went to the other side so it is negetive

OpenStudy (anonymous):

oh ok sin^2x=1-cos^2x

OpenStudy (anonymous):

yes now go back to your question and look if it is similar to this

OpenStudy (anonymous):

i dont understand how you decided to use that identity...

OpenStudy (anonymous):

and thats not the right answer..

OpenStudy (anonymous):

\[\cos x(\color{red}{1-\cos ^2x})=\cos x \color{red}{\sin^2x}\] red part from identity

OpenStudy (anonymous):

This isnt easy =..erg.. alright

OpenStudy (anonymous):

is this the answer or not

OpenStudy (anonymous):

no

OpenStudy (anonymous):

cos(x)sin^2(x) is the answer

OpenStudy (anonymous):

yes thats the one i wrote

OpenStudy (anonymous):

i dont see it...

OpenStudy (primeralph):

Still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please

OpenStudy (primeralph):